OpenStudy (anonymous):

Solve the equation. x^2+ 8x + 20=0

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

ok nvm what i just said the zeros are complex is that ok?

OpenStudy (anonymous):

yes I did. Basically I did, but I couldn't find what the factorials would be..

OpenStudy (anonymous):

east to complete the square, above answer is incorrect, since \((x+5)(x+3)=x^2+8x+15\)

OpenStudy (anonymous):

I just guessed (x+2) (x+10)=0 ... but i'm sure that's wrong.

OpenStudy (anonymous):

\[x^2+8x+20=0\] \[x^2+8x=-20\] \[(x+4)^2=-25+16=-4\] \[x+4=\pm\sqrt{-4}=\pm2i\]\[x=-4\pm 2i\]

OpenStudy (anonymous):

it does not have real zeros if you do not allow complex numbers in the class, then there are no solutions

OpenStudy (anonymous):

(x+4)2=−25+16=−4 how's you get this part?

OpenStudy (anonymous):

*how'd

OpenStudy (anonymous):

i completed the square did you get to that yet, or no?

OpenStudy (anonymous):

WE learned about it awhile ago, but I forget...

OpenStudy (anonymous):

and i made a mistake, the \(-25\) was a typo, it should have been \(-20+16=-4\) but the rest is right

OpenStudy (anonymous):

are you ok up until \[x^2+8x=-20?\]

OpenStudy (anonymous):

cause the rest is real real easy i can walk you through it if you like

OpenStudy (anonymous):

yes I understand the first part if you could walk me through the rest that would be great. :)

OpenStudy (anonymous):

ok sorry i stepped away ready?

OpenStudy (anonymous):

yes Iam and it's fine

OpenStudy (anonymous):

ok we start with \[x^2+8x=-20\] what is half of \(8\)?

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so we are going to replace the left hand side by \[(x+4)^2=-20+\] and we need to know what goes after the \(+\) sign on the right what is \(4^2\)

OpenStudy (anonymous):

16?

OpenStudy (anonymous):

yes, and that is what goes on the right so you turn \[x^2+8x=-20\] in to \[(x+4)^2=-20+16=-4\]

OpenStudy (anonymous):

ohh okay I understand now. Thankyou so much!!!

OpenStudy (anonymous):

the reason this works (not that you have do write this out) is that \[(x+4)^2=x^2+8x+16\] so addes \(16\) to the left and therefore you have to add \(16\) to the right to compensate

OpenStudy (anonymous):

* added

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

oh okaay